# Deep RL Needs Deep Behavior Analysis: Exploring Implicit Planning by Model-Free Agents in Open-Ended Environments

### v1

@@ 4.3 Emergence of Behavioral Competencies Over Training
To examine how agent behavior evolves during training, we analyzed performance metrics and behavioral patterns across episodes. We find that ForageWorld supports staged learning: agents shift from undirected exploration to structured, goal-aligned strategies—transitions only partially visible in standard reward curves.
These patterns suggest temporally organized skill acquisition. To test this, we tracked how different behaviors emerged and stabilized over the course of training. We used a panel of behavioral metrics to characterize the timeline of skill acquisition and identify which behaviors drove early versus late performance gains (Figure 4, Figure 9).
Figure 4: Distinct behavioral competencies emerge over training, with early gains in exploration followed by refinement of survival strategies. Training dynamics show staged acquisition of task-aligned skills. Early learning emphasizes spatial exploration and arena coverage, while later training refines predator avoidance, tool use, and resource regulation. Metrics include: spatial uncertainty (normalized by distance from origin); distance from origin (early vs. late); state occupancy entropy of agent position; angular orientation variance; predator field-of-view exposure; tool-making rate (1 = one tool crafted, 2 = both); and food/water satiation levels.
Early in training, PPO-RNN agents exhibited a “fishing” strategy—remaining stationary near the origin for extended periods. Around 20,000 PPO iterations, agents underwent an abrupt shift in behavior across multiple metrics. This transition included longer-distance travel during both early exploration and later revisitation, more strategic trade-offs between water and food collection, a sharp increase in tool-making frequency, and the emergence of gradual gains in predator defense. This shift marks the emergence of higher-level strategy layers in the training process.
To evaluate the contribution of specific components of the environment and model, we compared our PPO-RNN baseline to several variants with targeted objectives or capacities removed. One key ablation altered the agent’s field of view (FOV) to include only the forward grid rows, rather than the default Craftax FOV of a 9×11 rectangle centered on the agent. This change increased the need for information-seeking behavior and more closely resembled animal vision, which is predominantly front-facing. We found that front-FOV agents explored farther earlier in training and often engaged in more localized, dense search patterns (Figure 14, Figure 15).
@@ 4.4 Recurrent State Representations Support Memory and Planning
A growing body of work in RL has examined whether model-free RNN agents can perform implicit planning without the explicit world models typically used for planning tasks [65, 66, 67]. This question also arises in neuroscience, where insects like ants and bees show planning-like behavior despite lacking mammalian hippocampal-entorhinal map systems [68, 69, 70]. Our analyses show that DRL agents revisit patches based on multiple factors such as proximity, predator history, and position prediction error (Figure 3, Figure 7). An agent’s position relative to the origin can also be decoded from its current RNN state—up to 50–100 timesteps into the past and future—supporting its capacity for long-horizon memory and planning (Figure 5).
Figure 5: RNN states encode allocentric position and temporal structure, revealing emergent memory and planning capacity in model-free agents. (A) A single decoding model was trained per agent to test whether spatial information was encoded in allocentric (relative to origin) or egocentric (relative to agent) coordinates. (B) Late-training decoding performance for egocentric distance remained at chance level across models. (C) Allocentric position could be decoded above chance up to 50–100 timesteps into the past and future. Arena count was varied to show that models did not rely on arena-specific cues. (D) Allocentric decoding improved over training. Error bars are larger in (D) because decoding was limited to timesteps 1000–6000 per arena, due to shorter survival in early-training (10k epoch) agents. All plots use average displacement per timestep as a chance baseline. Error bars reflect 95% confidence intervals.
@@ 4.5 Generalization and Modularity in Recurrent Dynamics
We next examine whether the internal representations learned by DRL agents generalize across environments and exhibit modular substructure. These properties are important not only for performance, but also for interpretability—revealing which units encode what, and under what conditions. Decoding performance was moderately higher in pruned (sparse) networks than in fully connected ones (Figure 16), confirming more efficient information encoding in sparse architectures—consistent with recent findings [56, 57, 58]. Removing the auxiliary objective for position prediction eliminated above-chance decoding performance (Figure 17), suggesting that the associated task performance loss (Figure 2) stems from the model’s inability to encode position information using the PPO objective alone. These results suggest that the auxiliary objective improved both agent performance and the interpretability of its internal representations—a dual benefit revealed only through neural analysis. Past and future position encoding relied on overlapping but functionally modular subpopulations of RNN units, which diverged more clearly at longer time horizons (Figure 18). Notably, a single decoding model generalizes across episodes—predicting past and future position from RNN states recorded in many different arenas for the same agent. This cross-episode generalization likely reflects the consistent spatial structure of the environment: although arena configurations vary, the 96$\times$96 grid layout and global orientation are preserved, providing a stable allocentric reference frame.
In summary, model-free RNN agents trained in open-ended tasks can develop structured internal representations that support memory, planning, and spatial generalization—without relying on explicit world models. ForageWorld induces naturalistic behavioral and neural motifs in DRL agents, and our analysis framework enables these motifs to be detected and interpreted using behavior-first tools. These results highlight the importance of behavior-first analysis for revealing structure that remains hidden under standard reward-based evaluation.
@@ 5 Discussion
Our results support two central claims: (1) ForageWorld induces behavior–neural motifs relevant to planning and memory, and (2) our toolkit reveals these motifs in ways that simple performance metrics cannot.
@@
We also challenge the view that sophisticated planning and memory require explicit world models or symbolic memory. Despite being trained with model-free objectives, RNN-based agents in our study exhibit revisitation, uncertainty-driven exploration, and allocentric encoding—behaviors typically associated with model-based agents. While recent work suggests that temporal credit assignment and environment interaction can give rise to planning-like behavior in RNNs [65, 66, 67], our study provides the first systematic evidence that such behavior can emerge in complex, naturalistic environments without world models. These findings highlight how internal interpretability is shaped not just by training outcomes, but by design choices that constrain or promote structure.
@@
We also find that architectural choices such as recurrence, pruning, and auxiliary losses significantly influence both performance and interpretability. Pruning improved decoding accuracy (Figure 16) while leaving performance and behavior substantively unchanged (Figure 2); the auxiliary position prediction objective enhanced predator-related behavior and spatial encoding structure (Figure 4, Figure 11, Figure 17) while also improving performance (Figure 2). This win-win fits with observations from neuroscience—biological agents do not simply maximize reward, but also engage in latent path learning and other subgoals [71]. Thus, there are benefits for RL researchers to embracing this complexity of learning objectives and bridging the gaps between artificial and natural intelligence.
@@
These findings also speak to neuroscience research on small nervous systems. Insects, for example, navigate using structured internal cues [69, 72]. The fact that similar behaviors emerge in DRL agents with only a few hundred recurrent units suggests that such agents may offer useful models for evaluating the cognitive capacities of small nervous systems. Taken together with behavioral consistency across arenas, the ability of a single decoding model to generalize across episodes likely reflects a stable, compass-like encoding of space—mirroring strategies used by insects and other animals to maintain orientation using global cues such as magnetic fields or polarized light [73, 74, 75]. This encoding may help explain how agents reliably revisit resource patches across environments with differing local layouts.
@@
More broadly, we dispute the idea that DRL models are too complex for neuroscience-style analysis. The brain is at least as complex—and likely more so—than current DRL agents [38]. If neuroscience tools cannot explain DRL agents with full behavioral and neural access, their utility for understanding the brain is limited. Testing these tools in complex simulation environments also addresses practical challenges in neuroscience, including low statistical power, high data costs, and limited experimental control [76]. By generating high-volume, ecologically grounded datasets with joint neural-behavior logging, we provide a platform for evaluating brain analysis techniques. Our diagnostic framework (Table 1) offers a guide for interpreting agent behavior and internal dynamics across diverse RL tasks. We hope this framework serves both as a roadmap for future neuroAI work and as a foundation for more interpretable RL evaluation.
Our results also highlight an opportunity for future work: incorporating curriculum learning into ForageWorld and related environments [77, 78]. The staged emergence of behavioral competencies we observed (e.g., in Figure 4) suggests that ForageWorld naturally supports curriculum-based training regimes. Future studies could leverage this structure—and the motif-based analysis framework—to probe how learning history shapes internal representations, and how staged training affects generalization, revisitation strategies, and neural dynamics.
@@
The 96$\times$96 grid size limit in Craftax prevents us from evaluating how our results generalize to even larger arenas. A second limitation is that we could not log the precise locations of all trees, rocks, and lakes, which limited our ability to study landmark-based navigation in detail. A third limitation is that incorporating a wider range of RL architectures with full behavioral-neural logging is currently non-trivial in this environment, which is why we focused on PPO-based models with shared components. We hope future work will explore these directions. We believe this work is likely to have a positive societal impact by fostering closer collaboration between neuroscientists, cognitive scientists, ethologists, and RL researchers in the analysis of deep RL agents—advancing all fields through shared insight into the structure and function of learned behavior.
@@ 6 Acknowledgments
We thank Albert Lee and members of the Rajan Lab for the helpful comments and feedback during the course of the project.
@@ Appendix A ForageWorld: Task and Training Details
We expand here on the ForageWorld environment and agent training setup described in subsection 3.1 and subsection 3.2 of the main text. This includes key environment parameters, procedural generation constraints, agent sensory input, and logging infrastructure used for downstream analysis.
@@ A.1 Procedural Arena Generation
Each arena is procedurally generated at the start of an episode using a fixed-size 96$\times$96 grid. Spawn points for cows (food) and lakes (drink) are initialized in random patches using Perlin noise, providing episodic variability alongside naturalistic structure. Obstacles are added to break line-of-sight and create navigational constraints. The agent begins at the center of the arena.
@@ A.3 Sensory Input and State Representation
At each timestep, the agent receives a 9$\times$11 grid-centered egocentric view of the environment (see Figure 1), along with an inventory vector containing current health, satiation levels (food and water), fatigue, and any collected items. This full observation is encoded by a feedforward neural net layer before being passed to the recurrent network.
@@ A.5 Training Hyperparameters
Agents were trained using Proximal Policy Optimization (PPO) [79] with Generalized Advantage Estimation (GAE). Each agent trained for 3 billion timesteps, resulting in a variable number of episodes—typically in the hundreds of thousands—depending on competence and learning speed. Episodes were capped at 100,000 timesteps, though this cap was rarely reached in practice. The PPO rollout horizon was 64, with a minibatch size of 8192 and a learning rate of 0.00025. We used the Adam optimizer with default momentum parameters. Additional hyperparameters are provided in Table 3. Policy and value losses were weighted using standard PPO coefficients (see Table 3). For stability, gradient norms were clipped to 1.0.
@@ A.6 Model Architecture Details for PPO-RNN
Our primary architecture combined Proximal Policy Optimization (PPO) [79] with a recurrent neural network, using Generalized Advantage Estimation (GAE) for stability. PPO optimized a clipped surrogate objective that balanced advantage maximization with policy stability:
@@
To handle partial observability in the task, the agent uses a gated recurrent unit (GRU) [80] to build memory over time. At each timestep $t$, the GRU took the current observation $o_{t}$ and previous hidden state $h_{t-1}$, producing an updated hidden state $h_{t}$. Both the policy $\pi(a_{t}\mid o_{\leq t})$ and the value estimate $V(o_{\leq t})$ are computed from $h_{t}$ via separate output heads. This architecture is shared between the actor and critic networks. We used a single-layer GRU with 512 hidden units. GRUs offer a gating mechanism that supports long-range memory retention while being more computationally efficient than LSTMs. This advantage makes them especially well suited for tasks involving memory and interpretability. We followed the implementation provided in the Craftax benchmark [11], which demonstrated that memory-equipped PPO agents perform substantially better on partially observed, long-horizon tasks. In total, our standard configuration contains approximately 6.5 million trainable parameters.

### v2

@@ 4.3 Emergence of Behavioral Competencies Over Training
To examine how agent behavior evolves during training, we analyzed performance metrics and behavioral patterns across episodes. We find that ForageWorld supports staged learning: agents shift from undirected exploration to structured, goal-aligned strategies—transitions only partially visible in standard reward curves. These patterns suggest temporally organized skill acquisition. To test this, we tracked how different behaviors emerged and stabilized over the course of training. We used a panel of behavioral metrics to characterize the timeline of skill acquisition and identify which behaviors drove early versus late performance gains (Figure 4, Figure 9).
Figure 4: Distinct behavioral competencies emerge over training, with early gains in exploration followed by refinement of survival strategies. Training dynamics show staged acquisition of task-aligned skills. Early learning emphasizes spatial exploration and arena coverage, while later training refines predator response, tool use, and pathing patterns. Metrics include: spatial uncertainty (normalized by distance from origin); distance from origin (early vs. late, meaning before vs. after the first 1500 timesteps in an episode); state occupancy entropy of agent position; angular orientation variance across 250 timestep intervals; predator field-of-view exposure; tool-making rate (1 = one tool crafted, 2 = both); and food/water satiation levels. Error bars are 95% CI. These plots can be compared to the generally linear performance gains seen in survival times across training in Figure 2.
Early in training, PPO-RNN agents exhibited a “fishing” strategy—remaining stationary near the origin for extended periods. After approximately 20,000 training iterations, agents underwent an abrupt shift in behavior across multiple metrics. This transition included longer-distance travel during both early exploration and later revisitation, more strategic trade-offs between water and food collection, a sharp increase in tool-making frequency, and the emergence of gradual gains in predator defense. This shift marks the emergence of higher-level strategies in the training process.
To evaluate the contribution of specific components of the environment and model, we compared our PPO-RNN baseline to several variants with targeted objectives or capacities removed. One key ablation altered the agent’s field of view (FOV) to include only the forward grid rows, rather than the default Craftax FOV of a 9×11 rectangle centered on the agent. This change increased the need for information-seeking behavior and more closely resembled animal vision, which is predominantly front-facing. We found that front-FOV agents explored farther earlier in training and often engaged in more localized, dense search patterns (Figure 14, Figure 15). We also trained a different DRL objective, an off-policy RL architecture called parallelized Q-network (PQN) [gallici2024simplifying] (Appendix E). PQN-LSTM models performed comparably to the PPO-GRU baseline Figure 29, and with overall similar learning histories and pathing (Figure 31, Figure 30), but with the noticeable behavioral differences of much lower rates of predator killing (i.e. reducing predator HP to 0) (Figure 31). Thus, PPO-GRU seems to converge to a predator fighting strategy, while PQN-LSTM converges to predator evasion.
@@ 4.4 Recurrent State Representations Support Memory and Planning
A growing body of work in RL has examined whether model-free RNN agents can perform implicit planning without the explicit world models typically used for planning tasks [ni2021recurrent, hayes2022practical, chung2024predicting]. This question also arises in neuroscience, where insects like ants and bees show planning-like behavior despite lacking mammalian hippocampal-entorhinal map systems [wehner1981searching, chittka2022mind, freire2023absence]. Our analyses show that DRL agents revisit patches based on multiple factors such as proximity, predator history, and position prediction error (Figure 3, Figure 7). An agent’s position relative to the origin can also be decoded from its current RNN state—up to 50–100 timesteps into the past and future—supporting its capacity for long-horizon memory and planning (Figure 5). Furthermore, approximately 100/512 neurons were found to be position-sensitive at the single neuron level using generalized linear model (GLM) analysis. The neural activity response to position change (GLM coefficient magnitude) increased with radial distance from the agent’s origin as well, suggesting a distance accumulation circuit that modulates origin-revisitation behavior (Appendix D).
Figure 5: RNN states encode allocentric position and temporal structure, revealing emergent memory and planning capacity in model-free agents. (A) A single decoding model was trained per agent to test whether spatial information was encoded in allocentric (relative to origin) or egocentric (relative to agent) coordinates. (B) Late-training decoding performance for egocentric distance remained at chance level across models. (C) Allocentric position could be decoded above chance up to approximately 50–100 timesteps into the past and future depending on the run. Training arena count per decoder was varied to show that models did not rely on arena-specific cues. (D) Allocentric decoding improved over training. Error bars are larger in (D) because decoding was limited to timesteps 1000–6000 per arena, due to shorter survival in early-training (10k epoch) agents. All plots use average displacement per timestep as a chance baseline. Error bars reflect 95% CI.
@@ 4.5 Generalization and Modularity in Recurrent Dynamics
We next examine whether the internal representations learned by DRL agents generalize across environments and exhibit modular substructure. These properties are important not only for performance, but also for interpretability—revealing which units encode what, and under what conditions. Decoding performance was moderately higher in pruned (sparse) networks than in fully connected ones (Figure 16), confirming more efficient information encoding in sparse architectures—consistent with recent findings [frankle2019lottery, michel2019sixteen, tumma2023leveraging]. Removing the auxiliary objective for position prediction eliminated above-chance decoding performance (Figure 17), suggesting that the associated task performance loss (Figure 2) stems from the model’s inability to encode position information using the PPO objective alone. These results suggest that the auxiliary objective improved both agent performance and the interpretability of its internal representations—a dual benefit revealed only through neural analysis. Past and future position encoding relied on overlapping but functionally modular subpopulations of RNN units, which diverged more clearly at longer time horizons (Figure 18). Notably, a single decoding model generalizes across episodes—predicting past and future position from RNN states recorded in many different arenas for the same agent. This cross-episode generalization likely reflects the consistent spatial structure of the environment: although arena configurations vary, the $\times96$ grid layout and global orientation are preserved, providing a stable allocentric reference frame.
In summary, model-free RNN agents trained in open-ended tasks can develop structured internal representations that support memory, planning, and spatial generalization—without relying on explicit world models. ForageWorld induces naturalistic behavioral and neural motifs in DRL agents, and our analysis framework enables these motifs to be detected and interpreted using behavior-first tools.
@@ 5 Discussion
We provide two main findings: (1) ForageWorld induces behavior–neural motifs relevant to planning and memory, and (2) our toolkit reveals these motifs in ways that simple performance metrics cannot.
@@
We also help challenge the common view in neuroscience that sophisticated planning and memory require explicit world models, mammalian-like brains, or symbolic memory [chittka2022mind, clement2024latent], a topic that is of growing interest in computer science as well [bush2025interpreting, dedieu2025improving]. Despite being trained with model-free objectives, RNN-based agents in our study exhibit revisitation, uncertainty-driven exploration, and allocentric encoding—behaviors typically associated with model-based agents. While recent work suggests that temporal credit assignment and environment interaction can give rise to planning-like behavior in RNNs [ni2021recurrent, hayes2022practical, chung2024predicting], our study provides the first systematic evidence that such behavior can emerge in complex, naturalistic environments without world models. Thus, internal interpretability is shaped not just by training outcomes, but by design choices that constrain or promote structure.
@@
We also find that architectural choices such as recurrence, pruning, and auxiliary losses significantly influence both performance and interpretability. Pruning improved decoding accuracy (Figure 16) while leaving performance and behavior substantively unchanged (Figure 2); the auxiliary position prediction objective enhanced predator-related behavior and spatial encoding structure (Figure 4, Figure 11, Figure 17) while also improving performance (Figure 2). This win-win fits with observations from neuroscience—biological agents do not simply maximize reward, but also engage in latent path learning and other subgoals [clement2024latent]. Thus, there are benefits for RL researchers to embracing this complexity of learning objectives and bridging the gaps between artificial and natural intelligence.
@@
These findings also speak to neuroscience research on small nervous systems. Insects, for example, navigate using structured internal cues [chittka2022mind, couzin2025]. The fact that similar behaviors emerge in DRL agents with only a few hundred recurrent units suggests that such agents may offer useful models for evaluating the cognitive capacities of small nervous systems. Taken together with behavioral consistency across arenas, the ability of a single decoding model to generalize across episodes likely reflects a stable, compass-like encoding of space—mirroring strategies used by insects and other animals to maintain orientation using global cues such as magnetic fields or polarized light [wehner1976polarized, wiltschko2012magnetic, heinze2014polarized]. This encoding strategy may explain how DRL agents reliably plan navigation across environments with differing local layouts.
@@
More broadly, we dispute the idea that DRL models are too complex for neuroscience-style analysis. The brain is at least as complex—and likely more so—than current DRL agents [saxe2021if]. If neuroscience tools cannot explain DRL agents with full behavioral and neural access, their utility for understanding the brain is limited. Testing these tools in complex simulation environments also addresses practical challenges in naturalistic neuroscience, including low statistical power, high data costs, and limited experimental control [nastase2020keep]. By generating high-volume, ecologically grounded datasets with joint neural-behavior logging, we provide a platform for evaluating brain analysis techniques. Our diagnostic framework (Table 1) offers a guide for interpreting agent behavior and internal dynamics across diverse RL tasks. We hope this framework serves both as a roadmap for future neuroAI work and as a foundation for more interpretable RL evaluation.
@@
The $\times96$ grid size limit in Craftax prevents us from evaluating how our results generalize to even larger arenas. A second limitation is that we could not log the precise locations of all trees, rocks, and lakes, which limited our ability to study landmark-based navigation in detail. A third limitation is that incorporating a wider range of RL architectures with full behavioral-neural logging is currently non-trivial in this environment, which is why we focused on PPO- and PQN-based models with shared components. We believe this work is likely to have a positive societal impact by fostering closer collaboration between neuroscientists, cognitive scientists, ethologists, and RL researchers in the analysis of deep RL agents—advancing all fields through shared insight into the structure and function of learned behavior.
@@ 6 Acknowledgments
We thank Albert Lee and members of the Rajan Lab for the helpful comments and feedback during the course of the project. We also thank the research assistants Yang Wu and Yunxin Wu for their help during revisions. R.S., R.P.B. and K.R. were collectively funded by the NIH (RF1DA056403, U01NS136507), James S. McDonnell Foundation (220020466), Simons Foundation (Pilot Extension-00003332-02, McKnight Endowment Fund, CIFAR Azrieli Global Scholar Program, NSF (2046583), Harvard Medical School Neurobiology Lefler Small Grant Award, Harvard Medical School Dean’s Innovation Award, and Harvard Medical School Leonard and Isabelle Goldenson Fellowship. This work has been made possible in part by a gift from the Chan Zuckerberg Initiative Foundation to establish the Kempner Institute for the Study of Natural and Artificial Intelligence at Harvard University. F.B.B. received support from Fulbright Program, Jansons Legat, and the Norwegian State Educational Loan Fund.
7 Appendix
This Appendix provides additional details on the ForageWorld environment and agent training setup, supplementing the descriptions in subsection 3.1 and subsection 3.2 of the main text. It includes key environment parameters, procedural generation constraints, agent sensory input, and logging infrastructure used for downstream analysis. We also provide supplementary analyses and figures that expand on the main results, including trajectory examples, revisitation models, training dynamics, ablations, and neural decoding structure.
@@ Appendix A ForageWorld: Task and Training Details
We designed ForageWorld to balance between task complexity, transparency to analysis, training efficiency (using JAX GPU acceleration), biological plausibility in navigational planning computations, and connection to joint RL and recurrent memory architectures that are of particular interest to neuroscientists. More complex deep RL tasks exist, but many such tasks (for example, Minecraft) are much more difficult to study and in particular more difficult to instrument for behavior analysis. In addition, the increased computational cost of experiments in such environments is prohibitive to some statistical analysis, and many computational neuroscientists in particular do not have access to compute resources sufficient to run experiments in such complex environments. Relative to common neuroscience tasks used to study foraging, e.g. n-armed bandit tasks which lack spatial or temporal components, ForageWorld pushes the complexity boundary on the neuro-side of neuroAI while still remaining tractable for neuroscience tools and neuroscientists to study. Further, ForageWorld as it exists is not trivial to learn— our agents require billions of timesteps of experience to train to competency on the task, and Craftax, the benchmark which we build upon to develop ForageWorld, has not been solved by any agent or architecture to date.
In the following section, we expand here on the ForageWorld environment and agent training setup described in subsection 3.1 and subsection 3.2 of the main text. This includes key environment parameters, procedural generation constraints, agent sensory input, and logging infrastructure used for downstream analysis.
@@ A.1 Procedural Arena Generation
Each arena is procedurally generated at the start of an episode using a fixed-size $\times96$ grid. Spawn points for cows (food) and lakes (drink) are initialized in random patches using Perlin noise, providing episodic variability alongside naturalistic structure. Obstacles are added to break line-of-sight and create navigational constraints. The agent begins at the center of the arena.
@@ A.3 Sensory Input and State Representation
[…]
@@ A.5 Training Hyperparameters
Agents were trained using Proximal Policy Optimization (PPO) [schulman2017proximal] with Generalized Advantage Estimation (GAE). Each agent trained for 3 billion timesteps, resulting in a variable number of episodes—typically in the hundreds of thousands—depending on competence and learning speed. Episodes were capped at 100,000 timesteps, though this cap was rarely reached in practice. The PPO rollout horizon was 64, with a minibatch size of 8192 and a learning rate of 0.00025. We used the Adam optimizer with default momentum parameters. Additional hyperparameters are provided in Table 3. Policy and value losses were weighted using standard PPO coefficients (see Table 3). For stability, gradient norms were clipped to 1.0.
@@ A.6 Model Architecture Details for PPO-RNN
Our primary architecture combined Proximal Policy Optimization (PPO) [schulman2017proximal] with a recurrent neural network, using Generalized Advantage Estimation (GAE) for stability. PPO optimized a clipped surrogate objective that balanced advantage maximization with policy stability:
@@
To handle partial observability in the task, the agent uses a gated recurrent unit (GRU) [chung2014empirical] to build memory over time. At each timestep $t$, the GRU took the current observation $o_{t}$ and previous hidden state $h_{t-1}$, producing an updated hidden state $h_{t}$. Both the policy $\pi(a_{t}\mid o_{\leq t})$ and the value estimate $V(o_{\leq t})$ are computed from $h_{t}$ via separate output heads. This architecture is shared between the actor and critic networks. We used a single-layer GRU with 512 hidden units. GRUs offer a gating mechanism that supports long-range memory retention while being more computationally efficient than LSTMs. This advantage makes them especially well suited for tasks involving memory and interpretability. We followed the implementation provided in the Craftax benchmark [matthews2024craftax], which demonstrated that memory-equipped PPO agents perform substantially better on partially observed, long-horizon tasks. In total, our standard configuration contains approximately 6.5 million trainable parameters.
@@
PQN-specific Parameters
$\epsilon$ Start 1.0 PQN default
$\epsilon$ Finish 0.05 Parameter sweep of {0.005, 0.01, 0.05, 0.1}
$\epsilon$ Decay 1.0 Parameter sweep of {0.1, 1.0, 2.0}
Total training timesteps 12,000,000,000 Chosen to allow PQN time to explore+converge
Total decay timesteps 6,000,000,000 Parameter sweep of {1B, 6B, 12B}
Num environment steps per update 128 PQN default
$\lambda$ 0.5 PQN default
@@
A.7 Future directions in ForageWorld
The Craftax platform also provides an opportunity for future work not yet explored here: incorporating curriculum learning into ForageWorld and related environments [leibo2019autocurricula, jiang2023minimax]. The staged emergence of behavioral competencies we observed (e.g., in Figure 4) suggests that ForageWorld naturally supports curriculum-based training regimes and Craftax is compatible with JAX curriculum learning packages. We hope future work will explore curriculum learning [leibo2019autocurricula, jiang2023minimax]. Future studies could leverage this structure—and the motif-based analysis framework—to probe how learning history shapes internal representations, and how staged training affects generalization, revisitation strategies, and neural dynamics. Along this direction, it would also be interesting to test LLM models’ foraging behavior, as navigation is an underexplored topic in LLM literature generally [qiao2025open] (though note that transformers and LLMs would be harder to analyze than the simpler RNNs chosen here for their brain-like recurrent activity). However, Craftax is currently not coded for language-based labels of state variables so additional modification would have to be done to support LLM tests.
Lastly, note that ForageWorld does not support multi-agent environments yet, but Craftax recently supports multi-agent designs [omari2025multi].